# Giuseppe Scanniello+

### v1

@@ IV-A Research Question
The primary objective of the experimental assessment is to assess preliminarily the validity of the developed predictive model for the stratification of patients at HR risk. In particular, we aimed to show whether the proposed predictive model, which is based on two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models, outperforms traditional ML models trained on the entire feature set (i.e., the features used in the two specialized models). To that end, we considered three baseline models–i.e., Decision Tree, Random Forest, and Support Vector Regression (SVR)– and compared their results (e.g., accuracy, precision, and sensitivity) with those of our prediction model. We choose these baselines because they represent the leading ML techniques in several AI application contexts and medical domains [16], in particular.
  • Decision Tree. It is a tree-structured model that recursively splits the data based on feature values to predict an outcome. It is intuitive and interpretable and can handle both classification and regression tasks.
  • Random Forest. It is an ensemble learning method that builds multiple decision trees on random subsets of data and features and aggregates their predictions. It improves generalization by reducing overfitting (e.g., as compared with the Decision tree) and increasing robustness. It is well-suited for handling high-dimensional data and capturing complex patterns.
  • Support Vector Regression. It is a regression variant of Support Vector Machines that aims to find a function within a specified error margin (epsilon) while maximizing model flatness. It captures non-linear relationships using kernel functions.
For the baseline models, as we did for our predictive model, the individuals are arranged into the two groups according to the lifespan prediction, i.e., if it is less than 440 days, the individual is considered at risk, otherwise not.
According to the primary objective of our experimental assessment, we defined and investigated the following main Research Question:
RQ: Does our predictive model outperform leading ML techniques in classifying at-risk patients?
@@ IV-B Experimental Methodology
To study RQ, we conducted a comparative analysis between our model and the three baselines mentioned before. These baselines, differently from our model, were trained using the entire set of 33 selected features in Table I. All models were trained and evaluated using the final dataset, consisting of 182 patient records, with 20% of the data reserved for testing. Baselines were implemented using standard algorithms from the scikit-learn library in Python. Just as for training the models underlying our meta-model, the baseline models have been optimized using a five-fold cross-validation strategy and GridSearchCV, which systematically tests various combinations of hyperparameters. This methodology ensures that all models were evaluated using the same data splits, enabling a fair comparison when studying RQ.
@@ IV-C Metrics to Assess the Predictions of the Models
The evaluation metrics used are standard and widely accepted in the application of ML techniques within the medical domain [17]. The used metrics are:
  • Accuracy: It represents the proportion of patients that are correctly identified and it is computed as follows:
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}\%$
    This metric assumes values between 0% and 100%. The higher the value, the better it is in terms of the accuracy of the prediction. For the sake of clarity and completeness, TPs (True Positives) refer to patients correctly identified as people at HF risk, while TNs (True Negatives) refer to patients correctly identified as patients not at HF risk. Instead, FPs (False Positives) represent patients mistakenly identified as at risk, while FNs (False Negatives) represent patients mistakenly identified as not at risk.
  • Precision (for the at-risk class): It indicates the proportion of patients predicted as at risk who truly are at risk. It is calculated as follows:
    $\text{Precision}=\frac{TP}{TP+FP}\%$
    Also, the precision metric assumes values between 0% and 100%. The higher the value, the better the correctness of the prediction is.
  • Sensitivity (Recall for the at-risk class): It measures the model’s ability to correctly identify patients at risk and is defined as follows:
    $\text{Sensitivity}=\frac{TP}{TP+FN}\%$
    Also, this metric assumes values between 0% and 100%. The higher the value, the better the sensitivity of the model is.
  • F1-Score (for the at-risk class): It is the balanced harmonic mean of precision and sensitivity. It offers a balanced measure when the costs of false positives and false negatives differ and is computed as follows:
    $F1\text{-Score}=2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}\%$
    This metric assumes values between 0% and 100%. The higher the value, the better it is.
  • Diagnostic Odds Ratio (DOR): It is a comprehensive metric that integrates sensitivity and specificity. It is computed as follows:
    $DOR=\frac{TP\times TN}{FP\times FN}$
    The higher the value for DOR, the better it is.
It is important to highlight that, since the model is intended to stratify patients by identifying those at risk for inclusion in the telemonitoring program, minimizing FN is a primary objective. In this context, high precision is particularly desirable, as it ensures that most of the patients flagged as “at risk” truly need intervention. Nevertheless, achieving high values across all evaluation metrics remains a desirable goal to ensure overall model reliability.
TABLE II: Medical evaluation metrics of the proposed model
Metric Value Explanation of the Results
Accuracy 84% 84% of patients are correctly identified as at risk.
Precision (At-Risk) 79% 79% of patients flagged as at risk truly are at risk.
Sensitivity (At-Risk) 95% 95% of at-risk patients are correctly identified.
F1-Score (At-Risk) 86% Good balance between precision and sensitivity.
DOR 45.6 Strong discriminative ability.
@@ V Results
In this section, we highlight the main findings. We conclude by showing possible practical implications and threats that could affect the validity of the obtained results.
@@ V-A Main Findings and Discussion
The performance results of our predictive model are summarized in Table II. While the overall accuracy of 84% may appear moderate at first glance, the high sensitivity (95%) ensures that nearly all high-risk patients are correctly identified—an essential requirement for patient monitoring in the PrediHealth project. The precision of 79% reflects a relatively high number of false positives, which is acceptable given the clinical context and the overall objective of the PrediHealth project by minimizing FNs. Furthermore, the high Diagnostic Odds Ratio (45.6) underscores the model’s strong discriminative power between high- and low-risk patients, reinforcing its suitability for deployment within PrediHealth.
The results in Table III suggest that the proposed model significantly outperforms the baseline machine learning models. It achieves an accuracy of 84%, compared to 57% for the Decision Tree, 60% for the Random Forest, and 62% for SVR. Notably, the sensitivity of our model reaches 95% which is well above the 60% (Decision Tree and Random Forest) and 50% (SVR) by suggesting a strong ability to correctly identify high-risk patients as compared with the baseline models. The precision of our model (79%) is slightly higher than that of SVR (71%), which has the highest precision among the baseline models. This indicates that our model has a better balance between minimizing false positives and ensuring the inclusion of at-risk patients in the telemonitoring program.
Additionally, the proposed model achieves an F1-score of 86%, outperforming Decision Tree 60% and Random Forest (62%), and Support Vector Regression (59%). Its DOR (45.6) further confirms its superior ability to distinguish between high-risk and low-risk patients, compared to Decision Tree (1.69), Random Forest (2.14), and SVR (3.25). These improvements highlight the model’s reliability for HF risk stratification of patients.
@@
Decision Tree Accuracy 57%
Precision (At-Risk) 60%
Sensitivity (At-Risk) 60%
F1-Score (At-Risk) 60%
DOR 1.69
Random Forest Accuracy 60%
Precision (At-Risk) 63%
Sensitivity (At-Risk) 60%
F1-Score (At-Risk) 62 %
DOR 2.14
Support Vector Regression Accuracy 62%
Precision (At-Risk) 71%
Sensitivity (At-Risk) 50%
F1-Score (At-Risk) 59%
DOR 3.25
@@ V-B Implications
Here, we focus on the implications of our results from the perspectives of practitioners and researchers, setting aside their specific relevance to the PrediHealth project. This choice reflects our view that the project’s implications are already evident, and it allows us to emphasize that our findings hold significance beyond the boundaries of the project itself. Specifically, we developed a predictive model for HF risk stratification of patients using a modified stacking ensemble approach applied to clinical and echocardiographic data. It achieved high sensitivity while maintaining an adequate accuracy, using a 440-day lifespan threshold. This outcome suggests that a modified stacking ensemble approach can effectively identify a significant portion of high-risk HF patients. This is relevant for practitioners in the healthcare field, as the model functions as a valuable screening tool. Its high sensitivity supports its role in initial patient stratification, prompting closer monitoring. Researchers could be interested in improving our model in order to enhance its accuracy while maintaining high sensitivity, potentially by integrating additional data modalities or optimizing our stacking ensemble approach.
Our overall model, which separates clinical and echocardiographic features into distinct specialized models before integrating their predictions via a meta-model, presents a potential advantage, as our initial results suggest. Similar to how modularity improves software systems, this structure may enhance the interpretability of risk scores for clinicians. Practitioners could benefit from this increased transparency, as it allows them to discern whether a patient’s risk is primarily driven by general clinical status or specific cardiac parameters derived from echocardiography. Researchers should empirically validate the interpretability and robustness of this architecture, comparing it to monolithic models trained on all features flatly. Our initial results seem to justify this future research direction and also to explore optimal meta-learning strategies for integrating predictions from specialized models by assessing the scalability when incorporating additional features.
The predictive model was trained and evaluated on real patient data/records, which underwent extensive pre-processing, including feature selection and the removal of records with missing data, ultimately resulting in a dataset of 182 instances. Researchers could validate the model’s performance and generalizability on larger, multi-center datasets representing diverse patient populations. A particular focus on HFpEF cases, as discussed in our related work, would be valuable. Researchers could also be interested in investigating the impact of different missing data imputation techniques and the development of automated, context-aware feature selection methods to enhance the model’s robustness and reliability.
@@ V-C1 External validity
Our predictive model was developed and evaluated using 182 samples. Generalizing outcomes from this clinically specific cohort poses a threat to external validity. The model’s performances might differ in other populations with varying demographic profiles, healthcare practices, or different diagnoses. Furthermore, the specific set of 33 selected features might not be universally collected or hold the same predictive power in different clinical contexts. Although our study provides initial evidence for the feasibility of this model within the PrediHealth project, caution is needed when generalizing our outcomes.
@@ V-C2 Internal Validity
Internal validity threats mostly concern factors internal to our study design that might have influenced the results. The main internal validity threat of the study concerns the limited number of baselines used for comparison. This is because this does not exclude that there could be other ML models with better performance. On the other hand, the models used as comparison baselines are among the most used in the medical field, justifying their choice.
@@ V-C3 Construct Validity
As for this kind of threats to validity, one key concern is how we define “HF risk” as survival of fewer than 440 days after characterization. While this provides a clear target, it may not fully capture the broader clinical view of when a patient needs intervention or faces a high risk of worsening health, which can lead to serious complications beyond death. As for the metrics used to assess our predictive model and baselines, we used standard performance measures such as sensitivity and accuracy. While these provide useful information, they possibly do not give a complete picture. Lastly, the model assumes that medical records are accurate. Any errors in these records could reduce the reliability of predictions and limit its usefulness in real-world settings.
@@ V-C4 Conclusion validity
Our definition of the binary outcome (“at risk” vs. “not at risk”), chosen for technical reasons (dataset balance) and clinical relevance, could have affected outcomes. Using a different threshold could lead to different conclusions. Another internal validity threat concerns the choice of 33 features, which was based on clinical experience and data availability. This selection might leave out important predictors or include less useful ones, making it harder to fully represent the complex nature of HF risk.
@@ V-C5 Reliability validity
Reliability validity threats concern the possibility of replicating our study. Due to the sensitive nature of patient health information and regulations like GDPR, the raw patient-level dataset used in this study cannot be made publicly available. This inherently limits direct replication. To mitigate this, we have detailed the features used in the model and the design of the empirical study to assess its validity.
VI Conclusions
In this paper, we propose a predictive model based on machine learning (ML) techniques to identify patients at risk of HF. The model adopts an ensemble learning approach using a modified stacking technique, which combines the outputs of two specialized models—one based on clinical features and the other on echocardiographic features—via a meta-model. We evaluate the model on a real-world dataset, and the results demonstrate strong performance in stratifying patients by HF risk. In particular, the model achieves high sensitivity (95%), ensuring that nearly all high-risk patients are correctly identified. The overall accuracy is 84%, which, while moderate in general ML contexts, is acceptable given our priority of minimizing false negatives (patients mistakenly identified as at risk). This is especially relevant for the telemonitoring program of the PrediHealth research project. Indeed, our preliminary findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools—not only within the PrediHealth initiative but also more broadly in clinical practice—to facilitate early intervention and personalized patient management. We also compared our model against three baselines that treat features in a flat manner (i.e., without grouping them into clinical and echocardiographic categories). The results indicate that our model consistently outperforms these baselines, highlighting the benefit of feature specialization and structured integration.
Future research will focus on expanding the dataset to further validate the model, optimizing its interpretability for clinical applications, and integrating additional physiological and behavioral parameters to improve prediction. It is also worth noting that the current model was trained on real-world patient data from the Tuscany region of Italy. Due to the limited dataset size and its regional scope, the findings should be interpreted with caution; however, the promising initial results motivate conducting broader studies.
@@ Acknowledgment
This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001.

### v3

@@ IV-A Research Question
The primary objective of our initial experimental assessment is to validate the meta-model for the stratification of patients at HF risk. In particular, we aimed to show whether it outperforms traditional ML models trained on the entire feature set. According to our primary objective, we defined and studied the following main Research Question:
RQ
Does our predictive model outperform leading ML techniques in classifying at-risk patients?
In addition to studying the three ML-based models used in our meta-model, we considered two baselines–i.e., Decision Tree [^3] It is a tree-structured model that recursively splits the data based on feature values to predict an outcome. It is intuitive and interpretable and can handle both classification and regression tasks [20]. and Support Vector Classifier [^4] It is a classification variant of Support Vector Machines [21] designed to identify the optimal hyperplane that separates data points of different classes with the maximum margin. It captures non-linear relationships using kernel functions. (SVC). We choose these baselines because they represent the standard for comparison in several AI application contexts and medical domains [22, 23, 24].
@@ IV-B Experimental Methodology
To study RQ, we conducted a comparative analysis between our model and the baselines that were trained on the entire set of 33 features in Table I. All models considered in the experimental assessment were trained and evaluated using the final dataset with 20% of the data reserved for testing. Baselines were implemented using standard algorithms from the scikit-learn library in Python. Just as for training the models underlying our meta-model, the baseline models have been optimized using a five-fold cross-validation strategy and GridSearchCV, which systematically tests various combinations of hyperparameters. This methodology ensures that all models were evaluated using the same data splits, enabling a fair comparison when studying RQ.
@@ IV-C Metrics to Assess the Predictions of the Models
The evaluation metrics used are standard and widely accepted in the application of ML techniques within the medical domain [25]:
  • Accuracy: It represents the proportion of patients that are correctly identified, and it is computed as: $\frac{TP+TN}{TP+TN+FP+FN}\%$. This metric assumes values between 0% and 100%. The higher the value, the better it is in terms of the accuracy of the prediction. For the sake of clarity and completeness, TPs (True Positives) refer to patients correctly identified as people at HF risk, while TNs (True Negatives) refer to patients correctly identified as patients not at HF risk. Instead, FPs (False Positives) represent patients mistakenly identified as at risk, while FNs (False Negatives) represent patients mistakenly identified as not at risk.
  • Precision: It indicates the proportion of patients predicted as at risk who truly are at risk and is calculated as: $\frac{TP}{TP+FP}\%$. Also, precision assumes values between 0% and 100%. The higher the value, the better it is.
  • Sensitivity (or recall): It measures the proportion of actual positive patients that are correctly identified as having HF risk. It is computed as follows $\frac{TP}{TP+FN}\%$. Also, this metric assumes values between 0% and 100%. The higher the value, the better it is.
  • F1-Score: It is the balanced harmonic mean of precision and sensitivity and is computed as: $2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}\%$. This metric assumes values between 0% and 100%. The higher the value, the better it is.
  • Diagnostic Odds Ratio (DOR): It is a comprehensive metric that integrates sensitivity and specificity. It is computed as follows: $\frac{TP\times TN}{FP\times FN}$. The higher the value, the better it is.
It is important to highlight that since the model is intended to stratify patients by identifying those at risk for inclusion in the telemonitoring program, minimizing FN is a primary objective. In this context, high precision is particularly desirable, as it ensures that most of the patients at risk truly need to take part in the telemonitoring program. Nevertheless, achieving high values across all evaluation metrics remains a desirable goal to ensure overall model reliability.
TABLE II: Medical evaluation metrics of the proposed model and specialized ones
Model Metric Value
Meta-model Accuracy 78%
Precision 70%
Sensitivity 91%
F1-Score 79%
DOR 20
Clinical model Accuracy 72%
Precision 67%
Sensitivity 79%
F1-Score 78%
DOR 16
Echocardiographic model Accuracy 78%
Precision 71%
Sensitivity 88%
F1-Score 78%
DOR 16.31
Random Forest Classifier Accuracy 76%
Precision 69%
Sensitivity 88%
F1-Score 76%
DOR 14.5
@@ V Results
In this section, we highlight the main findings. We conclude by showing possible practical implications and threats that could affect the validity of the results obtained.
@@ V-A Main Findings and Discussion
The results of our meta-model are summarized in Table II (top). Although the overall precision of 78% appears moderate at first glance, the high sensitivity (91%) ensures that nearly all patients at HF risk are correctly identified - an essential requirement for patient monitoring in the PrediHealth project. The precision of 70% reflects a relatively high number of false positives, which is acceptable given the overall objective of the PrediHealth project by minimizing FNs. Furthermore, the high DOR (20) underscores the model’s strong discriminative power between patients at HF risk and those not. This reinforces the suitability of our meta-model for its deployment within the PrediHealth project. In Table II, we also report the results for each model underlying our meta-model. We can note that the meta-model allows obtaining better results as compared with its underlying ML-based models. The only difference concerns the precision value of the Echocardiographic model, which is 71%, while the precision value is 70% for the meta-model. These results support the conclusion that combining the predictions of the three underlying models contributes to improved performance, indicating that each model captures complementary aspects of the data.
@@
Decision Tree Accuracy 67%
Precision 60%
Sensitivity 79%
F1-Score 68%
DOR 4.81
SVC Accuracy 68%
Precision 63%
Sensitivity 72%
F1-Score 68%
DOR 4.76
The results in Table III suggest that the meta-model outperforms the baseline models. It achieves an accuracy of 78%, compared to 67% for the Decision Tree and 68% for SVC. The sensitivity of our model reaches 91% which is well above the 79% and 72% for Decision Tree and SVC, respectively. This suggests a better capability of the meta-model to correctly identify patients at HF risk and not if compared with the baselines. The precision of the meta-model (70%) is higher than the precision of the baselines (60% and 63% for Decision Tree and SVC, respectively).. This indicates that the meta-model has a better balance between minimizing FPs and ensuring the inclusion of at-risk patients in the telemonitoring program of PrediHealth. Additionally, the proposed model achieves an F1-score of 79%, outperforming Random Forest and SVC (68% for both models). Its DOR (20) further confirms its superior ability to distinguish between patients at HF risk and those not, compared to Decision Tree (4.81) and SVC (4.76). These improvements highlight the reliability of the model for the aim for which it is developed.
@@ V-B Implications
Here, we focus on the implications of our results from the perspectives of practitioners and researchers, setting aside their specific relevance to the PrediHealth project. This choice reflects our view that the project’s implications are already evident, and it allows us to emphasize that our findings hold significance beyond the boundaries of the project itself. Specifically, we developed a predictive model using a modified stacking ensemble approach applied to clinical and echocardiographic data. It achieved high sensitivity while maintaining adequate accuracy. This outcome suggests that a modified stacking ensemble approach can effectively identify a significant portion of patients at HF risk. This is relevant for practitioners in the healthcare field, as the model works as a valuable screening tool. Its high sensitivity supports its role in initial patient stratification. Researchers could be interested in improving our model to enhance its accuracy while maintaining high sensitivity, potentially by integrating additional data modalities or optimizing the used stacking ensemble approach.
Our meta-model, which uses clinical and echocardiographic features and all the features flattened, presents a potential advantage, as our initial results suggest. Similar to how modularity improves software systems, this structure may enhance the interpretability of risk scores for practitioners. They could benefit from this increased transparency, as it allows them to discern whether a patient’s risk is primarily driven by general clinical status or specific cardiac parameters derived from echocardiography. This aspect is also relevant for researchers. They could also be interested in empirically validating the interpretability and robustness of this architecture. For example, we provided initial evidence by showing that combining the predictions of three ML-based models by a meta-model improves the final predictions. Therefore, our empirical assessment seems to justify future research on the use of the ensemble learning approach and to explore optimal meta-learning strategies for integrating predictions from different models.
The predictive model was trained and evaluated on real patient data/records, which underwent pre-processing, including feature selection and the removal of records with missing data. Researchers could validate the performance and generalizability of the model on larger multi-center datasets representing diverse patient populations. A particular focus on HFpEF cases, as discussed in our related work, would be valuable. Researchers could also be interested in investigating the impact of different missing data imputation techniques and the development of automated, context-aware feature selection methods to enhance the robustness and reliability of a predictive model.
@@ V-C1 External validity
Our predictive model was developed and evaluated using 357 samples. Generalizing outcomes from this clinically specific cohort poses a threat to external validity. The model’s performances might differ in other populations with varying demographic profiles, healthcare practices, or different diagnoses. Furthermore, the specific set of 33 selected features might not be universally collected or hold the same predictive power in different clinical contexts. Although our study provides initial evidence for the feasibility of this model within the PrediHealth project, caution is needed when generalizing our outcomes.
@@ V-C2 Internal Validity
The main internal validity threat of the study concerns the number of baselines used in the empirical assessment. There could be other ML models outperforming our meta-model. On the other hand, the models used as comparison baselines are among the most used in the medical field, justifying their choice.
@@ V-C3 Construct Validity
As for this kind of threat to validity, one key concern is how we define patients at HF risk. It may not fully capture the broader clinical view of when a patient needs intervention or faces a high risk of worsening health, which can lead to serious complications beyond death. As for the metrics used to assess our predictive model and baselines, we used standard performance measures. While these provide useful information, they possibly do not give a complete picture. Lastly, our model assumes that medical records are accurate. Any errors in these records could reduce the reliability of predictions and limit their usefulness in real-world settings.
@@ V-C4 Conclusion validity
Our definition of the binary outcome (at risk vs. not at risk), chosen for technical reasons (dataset balance) and clinical relevance, could have affected outcomes. Using a different threshold could lead to different conclusions. Another internal validity threat concerns the choice of 33 features, which was based on clinical experience and data availability. This might leave out important features or include less useful ones.
@@ V-C5 Reliability validity
Reliability validity threats concern the possibility of replicating our study. Due to the sensitive nature of patient health information and regulations like GDPR, the raw patient-level dataset used in this study cannot be made publicly available. This inherently limits direct replication. We aim to make the source code implementing the data pre-processing pipeline and the described ensemble model architecture available upon reasonable request, facilitating conceptual replication and verification of our research.
VI Conclusion
In this paper, we propose a predictive model based on machine learning (ML) techniques to identify patients at HF risk. The model adopts an ensemble learning approach using a modified stacking technique, which combines the outputs of three underlying ML-based models. We evaluate our model on a real-world dataset, and the results demonstrate strong performance in stratifying patients at HF risk. In particular, the model achieves high sensitivity (91%), ensuring that nearly all patients are correctly identified as being at HF risk or not. The overall accuracy is 78%, which is acceptable given our priority of minimizing false negatives (patients mistakenly identified as at risk). This is especially relevant for the telemonitoring program of the research project PrediHealth, for which the proposed predictive model has been developed. We also compared our model against some ML-based models, and the results indicate that our model outperforms these models.
@@ Acknowledgment
This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001. This work was also partially supported by the project “DHEAL – COM- Digital Health Solutions in Community Medicine” under the Innovative Health Ecosystem (PNC) - National Recovery and Resilience Plan (NRRP) program funded by the Italian Ministry of Health.